# Learning to Segment for Vehicle Routing Problems

### v1

@@ 4 Learning to Segment (L2Seg)
Effective FSTA decomposition requires accurate identification of stable and unstable edges in solution $\mathcal{R}$. In this section, we introduce Learning to Segment (L2Seg), a neural framework for predicting unstable edges to guide FSTA. There are two natural paradigms: 1) Non-autoregressive (NAR) and 2) Autoregressive (AR) models. NAR models, e.g., heatmap, offer global predictions with an efficient single forward pass. However, they lack conditional modeling to accurately capture local dependencies. For example, when one edge is unstable, nearby edges often show instability, though not all are unstable. NAR models may fail to distinguish them and tend to mark all neighboring edges as unstable. On the other hand, AR models, e.g., attention-based models, can more natively capture local dependencies. Yet, they may miss the crucial global structure. For example, when unstable edges are distributed across distant regions, AR models may struggle to recognize and model these broader patterns. Our approach offers three variants as shown in Figure 2: non-autoregressive (L2Seg-NAR), autoregressive (L2Seg-AR), and a synergized combination of both (L2Seg-SYN).
@@ 4.1 Neural Architecture
Input Feature Design. We propose enhanced input features for L2Seg to better distinguish unstable and stable edges (see Appendix A.1 for intuitions). Key features include node angularity relative to the depot and node internality, where the latter measures the proportion of nearest nodes within the same route. We consider two edge types: edges in the current solution $\mathcal{R}$ and edges connecting each node to their k-nearest neighbors. Appendix B.1 provides a detailed feature description.
Encoder. Given node features $\mathbf{X}=(\mathbf{x_{0}},\mathbf{x_{1}},\dots)$ and edge features $\mathbf{E}=\{\mathbf{e}_{0,1},\mathbf{e}_{0,2},\dots\}$, we compute the initial node embedding as $\mathbf{h}^{\mathrm{init}}_{i}\!=\!\text{Concat}(\mathbf{h}^{\mathrm{MLP}}_{i},\mathbf{h}^{\mathrm{POS}}_{i})\in\mathbb{R}^{2d_{h}}$, where $\mathbf{h}^{\mathrm{MLP}}_{i}$ and $\mathbf{h}^{\mathrm{POS}}_{i}$ are obtained by passing $\mathbf{x_{i}}$ through a multilayer perceptron (MLP) and an absolute position encoder [47], respectively. Next, we process the embeddings using $L_{\mathrm{TFM}}$ Transformer layers [47] with masks to prevent computation between nodes in different routes: $\mathbf{h}^{\mathrm{TFM}}_{i}\!=\!\text{TFM}\left(\mathbf{h}^{\mathrm{init}}_{i}\right)\in\mathbb{R}^{d_{h}}$. This step encodes local structural information from the current solution. Finally, we compute the node embeddings $\mathbf{H}^{\mathrm{GNN}}\!=\!\{\mathbf{h}^{\mathrm{GNN}}_{i}\in\mathbb{R}^{d_{h}}\mid i=0,\dots,|V|\}$ leveraging the global graph information by using $L_{\mathrm{GNN}}$ layers of a Graph Attention Network (GAT) [48], where $\mathbf{H}^{\mathrm{GNN}}\!=\!\text{GNN}\left(\mathbf{H}^{\mathrm{TFM}},\mathbf{E}\right)$.
@@
Autoregressive Decoder. To model the interdependence of unstable edges, the autoregressive decoder generates the unstable edge set sequentially following a sequence $a=\{x_{\pi_{0}},x_{\pi_{1}},\dots\}$. Motivated by classical local search where $k$ removed unstable edges are linked via $k$ newly inserted edges between their endpoints [49, 28], this sequence is generated by alternating between a deletion stage that identifies unstable edges and an insertion stage that introduces a pseudo-edge to more accurately bridge to the next unstable edge, ending when a special token $x_{\mathrm{end}}$ is selected. Note that the insertion stage does not add new edges to the solution; it solely models the interdependence of consecutive unstable edges, reflecting the local search principle that removed edges are typically reconnectable via new insertions. Formally, denote the set of edges within the current solution as $E_{\mathcal{R}}$. The decoding alternates between: (1) Deletion ($t=2k$): Selects an unstable edge $e_{\pi_{2k},\pi_{2k+1}}\in E_{\mathcal{R}}$ based on a target node, which is either initialized at the first step (see Section 4.3) or the one obtained from the previous insertion step; one of the two edges connected to this node in the current solution is then selected as unstable (more than two candidates may exist if the node is the depot); and (2) Insertion ($t=2k+1$): Selects an new edge $e_{\pi_{2k+1},\pi_{2k+2}}\notin E_{\mathcal{R}}$ that links to the endpoint of the last unstable edge removed, exploring $O(|V|)$ potential candidates to serve as a bridge to the next unstable target node (next unstable region). From $a$, we then identify the set of removed edges as the unstable edges, i.e., $E_{\text{unstable}}=\{e_{\pi_{0},\pi_{1}},e_{\pi_{2},\pi_{3}},\dots\}$. Both stages employ two principal modules: Gated Recurrent Units (GRUs) [50] to encode sequence context, and multi-head attention (MHA) [47] for node selection. The GRU’s initial hidden state is the average of all node embeddings: $\mathbf{h}^{\text{hidden}}_{0}=\frac{1}{|V|}\sum_{i=0}^{|V|}\mathbf{h}^{\text{GNN}}_{i}$. At step $t$, the sequence embedding is updated by $\mathbf{h}^{\text{seq}}_{t}=\mathrm{GRU}\bigl{(}\mathbf{h}^{\text{hidden}}_{t-1},\,\mathbf{h}^{\text{GNN}}_{\pi_{t-1}}\bigr{)}$, and the context embedding is formed by concatenating the embeddings of the initial node, the previous node, and the new sequence embedding: $\mathbf{H}_{t}^{\text{context}}=\mathrm{Concat}\bigl{(}\mathbf{h}^{\text{GNN}}_{\pi_{0}},\,\mathbf{h}^{\text{GNN}}_{\pi_{t-1}},\,\mathbf{h}^{\text{seq}}_{t}\bigr{)}$.
@@
Inspired by the decoder design in LEHD [25], we use two distinct MHA modules with $L^{\mathrm{MHA}}$ layers, to decode $x_{\pi_{t}}$. Specifically, considering the size of the action space (at most 2 for deletion and $O(|V|)$ for insertion), we utilize a shallow decoder ($L_{\mathrm{delete}}^{\mathrm{MHA}}=1$) during the deletion stage and a deeper decoder ($L_{\mathrm{insert}}^{\mathrm{MHA}}=4$) during the insertion stage. Let $\mathbf{H}^{a}_{t}\subseteq\mathbf{H}^{\mathrm{GNN}}$ denote the set of available nodes at step $t$. During the insertion stage, we also incorporate an additional candidate $\mathbf{h}^{\mathrm{end}}=\alpha\mathbf{h}^{\mathrm{GNN}}_{\pi_{0}}\!+\!(1-\alpha)\,\frac{1}{|V|}\sum_{i=0}^{|V|}\mathbf{h}^{\mathrm{GNN}}_{i}$, where $\alpha$ is a learnable parameter, to indicate termination of decoding, providing the AR model with the flexibility to determine the number of unstable edges. Formally, the decoding at step $t$ is given as follows; note that the first 3 dimensions of $\mathbf{H}^{(L^{\mathrm{MHA}})}$ corresponds to the context embeddings $\mathbf{H}_{t}^{\text{context}}$ and hence are masked from selection:
@@
$\begin{split}\mathbf{H}^{(0)}&=\mathrm{Concat}\bigl{(}\mathbf{H}_{t}^{\mathrm{context}},\,\mathbf{H}^{a}_{t}\bigr{)},\mathbf{H}^{(l)}=\mathrm{MHA}\bigl{(}\mathbf{H}^{(l-1)}\bigr{)},\\ u_{i}&=\begin{cases}(W_{q}\mathbf{h}^{\mathrm{c}})^{T}W_{k}\mathbf{h}_{i}^{(L^{\mathrm{MHA}})}\,\!\!/\!\,\sqrt{d_{h}},&\text{if }i>3,\\ -\infty,&\text{otherwise},\end{cases}\end{split}$ (2)
@@ 4.2 Training
We employ iterative solvers as look-ahead heuristics to detect unstable edges. We then utilize imitation learning to train L2Seg models, so as to replicate the behavior of the look-ahead heuristics.
Dataset Construction. Let the edges indicated by a solution $\mathcal{R}$ be $E_{\mathcal{R}}$, and the nodes indicated by an edge set $E$ be $V_{E}$. Given $P$ with the current solution $\mathcal{R}$, we first employ an iterative solver $\mathcal{S}$ to refine $\mathcal{R}$ and obtain $\mathcal{R}_{+}$. We then collect the differing edges between $\mathcal{R}$ and $\mathcal{R}_{+}$ as $E_{\mathrm{diff}}\!=\!\bigl{(}E_{\mathcal{R}}\setminus E_{\mathcal{R}_{+}}\bigr{)}\!\cup\!\bigl{(}E_{\mathcal{R}_{+}}\setminus E_{\mathcal{R}}\bigr{)}$ (including both the deleted and newly inserted edges). Next, we identify the set of unstable nodes $V_{\mathrm{unstable}}\!=\!V_{E_{\mathrm{diff}}}$, i.e., the set of nodes that are end points to some edge in $E_{\text{diff}}$. We empirically observe that solution refinement typically takes place between two adjacent routes. For the NAR model, we construct a dataset with binary labels. Each problem-label pair consists of a decomposed problem containing two adjacent routes and binary labels indicating whether each node is unstable (1) or stable (0). Formally, a node $x$ is labeled 1 if $x\in V_{\text{unstable}}$.
To capture local dependencies among unstable edges, we construct AR labels as node sequences that preserve these relationships. Nodes without local dependencies are naturally excluded through connected component partitioning. We thus obtain the set of connected components $\mathcal{K}$ induced by $E_{\mathrm{diff}}$ and select those containing nodes from at most two routes, denoted by $\mathcal{K_{\mathrm{TR}}}$. For all $K\in\mathcal{K_{\mathrm{TR}}}$, let $R_{i}$ and $R_{j}$ be the two routes ($\forall x\in K,\!x\in R_{i}\cup R_{j}$). We then form a subproblem $P_{K}$ containing those two routes with solution $\mathcal{R}_{K}=\{R_{i},R_{j}\}$. For any $K\in\mathcal{K_{\mathrm{TR}}}$, we extract a sequence of nodes $y_{K}\!=\!\{x_{\pi_{0}},x_{\pi_{1}},\dots,x_{\pi_{m}},x_{\mathrm{end}}\}$ as the label for the AR model (alternatively deleting and inserting edges). With the corresponding problem $P_{K}$, we construct problem–label pairs with the corresponding $y_{K}$ for the AR model. Additional details are in Appendix C.5..
@@ 4.3 Inference
We describe the synergized inference that combines the benefits of global structural awareness from NAR with the local precision from AR, followed by two simpler variants using only NAR or AR.
@@
Autoregressive Prediction (L2Seg-AR). Unlike L2Seg-SYN, L2Seg-AR exclusively utilizes the AR model. Instead of using the NAR model to detect unstable nodes, it assumes all nodes may be unstable, hence applying the $K$-means algorithm on all nodes. It then selects the node closest to each cluster center as the initial node for AR-based decoding.
@@ 5 Experiment
Our decomposition-based FSTA and L2Seg excel on large-scale problems. In this section, we first evaluate how L2Seg-AR, L2Seg-NAR, and L2Seg-SYN accelerate various learning and non-learning iterative solvers on large-capacity CVRPs with long subtours. Next, we compare L2Seg against state-of-the-art baselines on standard benchmark CVRP and VRPTW instances. Finally, we provide in-depth analyses of our pipeline. Additional results and discussions are available in Appendix D. Our code, pre-trained models, and the dataset will be released on Github (see Appendix E).
Table 1: Performance comparisons of our proposed L2Seg-NAR, L2Seg-AR, and L2Seg-SYN when accelerating three backbone solvers, LKH-3, LNS, and L2D, on the large-capacity CVRP instances. We report the objective value, improvement gain (%), and the time. The gains (the higher the better) are w.r.t. the performance of each backbone solver. Time limits were set to be 150s for CVRP2k and 240s for CVRP5k, respectively.
@@
Backbone Solvers. We integrate L2Seg with three representative backbones: LKH-3 [7] (classic heuristic), LNS [9] (decomposition framework), and L2D [4] (learning-guided hybrid solvers) to demonstrates the broad applicability of L2Seg. Details on backbones are provided in Appendix C.1.
Baselines. We include state-of-the-art classic solvers (LKH-3 [7], HGS [8]), neural solvers (BQ [16], LEHD [25], ELG [15], ICAM [51], L2R [42], SIL [5]), and learning-based divide-and-conquer methods (GLOP [40], TAM [41], UDC [26], L2D [4]). We rerun LKH-3, LNS, and L2D, and report results from [26, 5] for other baselines using the same benchmarks. See Appendix C.2 for details.
Data Distribution. We generate all training and test instances following prior works by Zheng et al. [26] for CVRP and Solomon [52] for VRPTW. See Appendix C.3 for details. For Section 5.1, results are averaged over 100 large-scale CVRP test instances at 2k and 5k scales (capacities 500 and 1,000, respectively). For Section 5.2, we follow standard NCO benchmarks, reporting averaged results on 1k, 2k, and 5k test datasets with 1,000 CVRP and 100 VRPTW instances per scale.
Evaluation and Metric. For all methods, we impose time limits of 150s, 240s, and 300s for CVRP1k, CVRP2k, and CVRP5k, and 120s, 240s, and 600s for VRPTW1k, VRPTW2k and VRPTW5k. Note that each solver may finish a few seconds ($<$ 10s) beyond its limit. We set $\eta=0.6$ and $n_{\text{KMEANS}}=3$ for our L2Seg. Additional parameters are listed in Appendix C.4. We report travel cost and per-instance solve time for all cases, and report percentage improvements over backbone solvers in Section 5.1 and gaps to HGS (the best heuristic solvers) for both CVRP and VRPTW in Section 5.2.

### v2

@@ 4 Learning to Segment (L2Seg)
We introduce Learning to Segment (L2Seg), a neural framework for predicting unstable edges to guide FSTA. We consider two paradigms: 1) Non-autoregressive (NAR) and 2) Autoregressive (AR) models. NAR models offer global predictions with an efficient single forward pass. However, they lack conditional modeling to accurately capture local dependencies. For example, when one edge is unstable, nearby edges often show instability but not all, but NAR models may fail to distinguish them and mark all neighboring edges as unstable. On the other hand, AR models can more natively capture local dependencies. Yet, they may miss the crucial global structure. For example, when unstable edges are distributed across distant regions, AR models may struggle to recognize and model these broader patterns. Our approach offers three variants as shown in Figure 2: non-autoregressive (L2Seg-NAR), autoregressive (L2Seg-AR), and a synergized combination of both (L2Seg-SYN).
@@ 4.1 Neural Architecture
Input Feature Design. We propose enhanced input features for L2Seg to better distinguish unstable and stable edges (see Appendix B.1 for intuitions). Key features include node angularity relative to the depot and node internality, where the latter measures the proportion of nearest nodes within the same route. We consider two edge types: edges in the current solution $\mathcal{R}$ and edges connecting each node to their k-nearest neighbors. Appendix C.2 provides a detailed feature description.
Encoder. Given node features $\mathbf{X}=(\mathbf{x_{0}},\mathbf{x_{1}},\dots)$ and edge features $\mathbf{E}=\{\mathbf{e}_{0,1},\mathbf{e}_{0,2},\dots\}$, we compute the initial node embedding as $\mathbf{h}^{\mathrm{init}}_{i}\!=\!\text{Concat}(\mathbf{h}^{\mathrm{MLP}}_{i},\mathbf{h}^{\mathrm{POS}}_{i})\in\mathbb{R}^{2d_{h}}$, where $\mathbf{h}^{\mathrm{MLP}}_{i}$ and $\mathbf{h}^{\mathrm{POS}}_{i}$ are obtained by passing $\mathbf{x_{i}}$ through a multilayer perceptron (MLP) and an absolute position encoder  (Vaswani, 2017), respectively. Next, we process the embeddings using $L_{\mathrm{TFM}}$ Transformer layers  (Vaswani, 2017) with masks to prevent computation between nodes in different routes: $\mathbf{h}^{\mathrm{TFM}}_{i}\!=\!\text{TFM}\left(\mathbf{h}^{\mathrm{init}}_{i}\right)\in\mathbb{R}^{d_{h}}$. This step encodes local structural information from the current solution. Finally, we compute the node embeddings $\mathbf{H}^{\mathrm{GNN}}\!=\!\{\mathbf{h}^{\mathrm{GNN}}_{i}\in\mathbb{R}^{d_{h}}\mid i=0,\dots,|V|\}$ leveraging the global graph information by using $L_{\mathrm{GNN}}$ layers of a Graph Attention Network (GAT)  (Veličković et al., 2017), where $\mathbf{H}^{\mathrm{GNN}}\!=\!\text{GNN}\left(\mathbf{H}^{\mathrm{TFM}},\mathbf{E}\right)$.
@@
Autoregressive Decoder. The autoregressive decoder models unstable edge interdependence by generating them sequentially as $a=\{x_{\pi_{0}},x_{\pi_{1}},\dots\}$. Following classical local search where $k$ removed edges are reconnected via $k$ new insertions (Funke et al., 2005), the sequence alternates between deletion (identifying unstable edges) and insertion (introducing pseudo-edges that bridge to the next unstable edge), terminating at $x_{\mathrm{end}}$. Note that the “insertion" stage is designed to model dependencies between consecutive unstable edges rather than actually “insert" edges into the solution. Formally, denote the set of edges within the current solution as $E_{\mathcal{R}}$. The decoding alternates between: (1) Deletion ($t=2k$): Selects an unstable edge $e_{\pi_{2k},\pi_{2k+1}}\in E_{\mathcal{R}}$ based on a target node, which is either initialized at the first step (see Section 4.3) or the one obtained from the previous insertion step; one of the two edges connected to this node in the current solution is then selected as unstable (more than two candidates may exist if the node is the depot); and (2) Insertion ($t=2k+1$): Selects an new edge $e_{\pi_{2k+1},\pi_{2k+2}}\notin E_{\mathcal{R}}$ that links to the endpoint of the last unstable edge removed, exploring $O(|V|)$ potential candidates to serve as a bridge to the next unstable target node (next unstable region). From $a$, we then identify the set of removed edges as the unstable edges, i.e., $E_{\text{unstable}}=\{e_{\pi_{0},\pi_{1}},e_{\pi_{2},\pi_{3}},\dots\}$. Both stages employ two principal modules: Gated Recurrent Units (GRUs)  (Chung et al., 2014) to encode sequence context, and multi-head attention (MHA)  (Vaswani, 2017) for node selection. The GRU’s initial hidden state is the average of all node embeddings: $\mathbf{h}^{\text{hidden}}_{0}=\frac{1}{|V|}\sum_{i=0}^{|V|}\mathbf{h}^{\text{GNN}}_{i}$. At step $t$, the sequence embedding is updated by $\mathbf{h}^{\text{seq}}_{t}=\mathrm{GRU}\bigl(\mathbf{h}^{\text{hidden}}_{t-1},\,\mathbf{h}^{\text{GNN}}_{\pi_{t-1}}\bigr)$, and the context embedding is formed by concatenating the embeddings of the initial node, the previous node, and the new sequence embedding: $\mathbf{H}_{t}^{\text{context}}=\mathrm{Concat}\bigl(\mathbf{h}^{\text{GNN}}_{\pi_{0}},\,\mathbf{h}^{\text{GNN}}_{\pi_{t-1}},\,\mathbf{h}^{\text{seq}}_{t}\bigr)$.
Inspired by the decoder design in LEHD  (Luo et al., 2023), we use two distinct MHA modules with $L^{\mathrm{MHA}}$ layers, to decode $x_{\pi_{t}}$. Specifically, considering the size of the action space (at most 2 for deletion and $O(|V|)$ for insertion), we utilize a shallow decoder ($L_{\mathrm{delete}}^{\mathrm{MHA}}=1$) during the deletion
Figure 4: Training data construction for the AR model. Re-optimization reveals deleted edges (blue/green dashed) and inserted edges (red dashed) forming connected components (circles). For each component, depth-first search generates node sequences alternating between deletion and insertion stages, terminated by an end token as the AR model’s training label.
@@
stage and a deeper decoder ($L_{\mathrm{insert}}^{\mathrm{MHA}}=4$) during the insertion stage. Let $\mathbf{H}^{a}_{t}\subseteq\mathbf{H}^{\mathrm{GNN}}$ denote the set of available nodes at step $t$. During the insertion stage, we also incorporate an additional candidate $\mathbf{h}^{\mathrm{end}}=\alpha\mathbf{h}^{\mathrm{GNN}}_{\pi_{0}}\!+\!(1-\alpha)\,\frac{1}{|V|}\sum_{i=0}^{|V|}\mathbf{h}^{\mathrm{GNN}}_{i}$, where $\alpha$ is a learnable parameter, to indicate termination of decoding, providing the AR model flexibility to determine the number of unstable edges. Formally, the decoding at step $t$ is given as follows; note that the first 3 dimensions of $\mathbf{H}^{(L^{\mathrm{MHA}})}$ corresponds to context embeddings $\mathbf{H}_{t}^{\text{context}}$ and hence are masked from selection:
@@
$\begin{split}&\mathbf{H}^{(0)}=\mathrm{Concat}\bigl(\mathbf{H}_{t}^{\mathrm{context}},\,\mathbf{H}^{a}_{t}\bigr),\\ &\mathbf{H}^{(l)}=\mathrm{MHA}\bigl(\mathbf{H}^{(l-1)}\bigr),\\ &u_{i}=\begin{cases}(W_{q}\mathbf{h}^{\mathrm{c}})^{T}W_{k}\mathbf{h}_{i}^{(L^{\mathrm{MHA}})}\,\!\!/\!\,\sqrt{d_{h}},&\text{if }i>3,\\ -\infty,&\text{O.W.},\end{cases}\end{split}$ (2)
@@ 4.2 Training
We employ iterative solvers as look-ahead heuristics to detect unstable edges. We utilize imitation learning to train L2Seg models to replicate the behavior of the look-ahead heuristics.
Dataset Construction. Let the edges in $\mathcal{R}$ be $E_{\mathcal{R}}$, and nodes indicated by edge set $E$ be $V_{E}$. Given $P$ with current solution $\mathcal{R}$, we first employ an iterative solver $\mathcal{S}$ to refine $\mathcal{R}$ and obtain $\mathcal{R}_{+}$. We then collect differing edges as $\mathcal{R}$ and $\mathcal{R}_{+}$ as $E_{\mathrm{diff}}\!=\!\bigl(E_{\mathcal{R}}\setminus E_{\mathcal{R}_{+}}\bigr)\!\cup\!\bigl(E_{\mathcal{R}_{+}}\setminus E_{\mathcal{R}}\bigr)$ (including both the deleted and newly inserted edges). Next, we identify the set of unstable nodes $V_{\mathrm{unstable}}\!=\!V_{E_{\mathrm{diff}}}$, i.e., the set of nodes that are end points to some edge in $E_{\text{diff}}$. We empirically observe that solution refinement typically takes place between two adjacent routes. For the NAR model, we construct a dataset with binary labels. Each problem-label pair consists of a decomposed problem containing two adjacent routes and binary labels indicating whether each node is unstable (1) or stable (0). Formally, a node $x$ is labeled 1 if $x\in V_{\text{unstable}}$. For the AR model, we construct labels as node sequences preserving local dependencies among unstable edges. Nodes without local dependencies are naturally excluded through connected component partitioning. We obtain connected components $\mathcal{K}$ induced by $E_{\mathrm{diff}}$ and select those spanning at most two routes, denoted $\mathcal{K}_{\mathrm{TR}}$. For each $K\in\mathcal{K}_{\mathrm{TR}}$ containing nodes from routes $R_{i}$ and $R_{j}$, we form a subproblem $P_{K}$ with solution $\mathcal{R}_{K}=\{R_{i},R_{j}\}$. From each component $K$ (dashed circles in Figure 4), we extract a node sequence $y_{K}=\{x_{\pi_{0}},x_{\pi_{1}},\dots,x_{\pi_{m}},x_{\mathrm{end}}\}$ by alternating between edge deletion and insertion operations (shown in Figure 4, second row). These problem-label pairs $(P_{K},y_{K})$ constitute the AR model training data.
@@ 4.3 Inference
We describe the synergized inference that combines the benefits of global structural awareness from NAR with the local precision from AR, followed by two variants using only NAR or AR.
@@
Autoregressive Prediction (L2Seg-AR). L2Seg-AR exclusively uses the AR model. Instead of using the NAR model, it assumes all nodes may be unstable, applying the $K$-means algorithm on all nodes. It then selects the node closest to each cluster center as the initial node for AR-based decoding.
@@ 5 Experiment
Our decomposition-based FSTA and L2Seg excel on large-scale problems. In this section, we first evaluate how L2Seg‑AR, L2Seg‑NAR, and L2Seg‑SYN accelerate various learning and non-learning iterative solvers on large‑capacity CVRPs with long subtours. Next, we compare L2Seg against state-of-the-art baselines on standard benchmark CVRP and VRPTW instances. Finally, we provide in‑depth analyses of our pipeline. Additional results on CVRPLib benchmarks, clustered CVRP, heterogeneous-demand CVRP, a case study, and further discussions are presented in Appendix E.
Table 1: Performance comparisons of our proposed L2Seg-NAR, L2Seg-AR, and L2Seg-SYN when accelerating three backbone solvers, LKH-3, LNS, and L2D, on the large‑capacity CVRP instances. We report the objective value, improvement gain (%), and the time. The gains (the higher the better) are w.r.t. the performance of each backbone solver. Time limits were set to be 150s for CVRP2k and 240s for CVRP5k, respectively.
@@
Backbone Solvers. We apply L2Seg to three representative backbones: LKH-3 (Helsgaun, 2017) (classic heuristic), LNS (Shaw, 1998) (decomposition framework), and L2D (Li et al., 2021) (learning-guided hybrid solvers) to demonstrates the broad applicability. See Appendix D.1 for details.
Baselines. We include state-of-the-art classic solvers (LKH-3 (Helsgaun, 2017), HGS (Vidal, 2022)), neural solvers (BQ (Drakulic et al., 2023), LEHD (Luo et al., 2023), ELG (Gao et al., 2024), ICAM (Zhou et al., 2024), L2R (Zhou et al., 2025a), SIL (Luo et al., 2024)), and learning-based divide-and-conquer methods (GLOP (Ye et al., 2024), TAM (Hou et al., 2023), UDC (Zheng et al., 2024), L2D (Li et al., 2021), NDS (Hottung et al., 2025)). We rerun LKH-3, LNS, L2D, and NDS and report results from Zheng et al. (2024), Luo et al. (2024) for other baselines using the same benchmarks. See Appendix D.2 for baseline setup details and Appendix D.3 for L2Seg hyperparameters.
Data Distribution. We generate all training and test instances following prior works Zheng et al. (2024) for CVRP and Solomon (1987) for VRPTW. See Appendix D.4 for details. For Section 5.1, results are averaged over 100 large‑scale CVRP test instances at 2k and 5k scales (capacities 500 and 1,000, respectively). For Section 5.2, we follow standard NCO benchmarks, reporting averaged results on 1k, 2k, and 5k test datasets with 1,000 CVRP and 100 VRPTW instances per scale.
Evaluation and Metric. We impose time limits of 150s, 240s, and 300s for CVRP1k, 2k, and 5k, and 120s, 240s, and 600s for VRPTW1k, 2k and 5k, where each solver may finish a few seconds (<< 10s) beyond its limit. We set $\eta=0.6$ and $n_{\text{KMEANS}}=3$ for our L2Seg. We report averaged cost and per-instance solve time for all cases, and report percentage improvements over backbone in Section 5.1 and gaps to HGS (the best heuristic solvers) for both CVRP and VRPTW in Section 5.2.